If 0.5 mole of BaCl2 is mixed with 0.2 mole of Na33PO4.the max no of moles of Ba3(PO4)2.?

Grade:12

2 Answers

Naveen Kumar
9 years ago
Step-I:
First write the balenced equation for this reaction,
3BaCl2.........+.......2Na3PO4.........=Ba3(PO4)2....+6NaCl
0.5 mol.....................0.2 mol...............x mol
Step-II:
In this balenced equation we should know the limiting reagent(the reagent which will be consumed first or which is in less amount)
So to know the limiting reagent, divide the gven moles by their respective coefficient and then the lowest value will be the limiting reagent.
BaCl2=0.5/3=0.167
Na3PO4=0.2/2=0.1
here clearly we can see that Na3PO4 is the limiting reagent and hence the product will be formed till theNa3PO4 is consumed. rest of the BaCl2 will be remaining after the reaction.
Now from the reaction, we can see that 2 moles ofNa3PO4 will give 1 mole of Ba3(PO4)2
Hence 1 mol ofNa3PO4will give ½ mole ofBa3(PO4)2
so 0.2 mole ofNa3PO4 ….........1/2 * 0.2 molBa3(PO4)2=0.1 mol
Shriya Mehrotra
9 years ago
thanks sir.